5x-4/5x+4 =2/3
I need help with this please.

Answer by checkley75(3666)

5X-4X/5=2/3
(25X-4X)/5=2/3
21X/5=2/3 NOW CROSS MULTIPLY
21X*3=5*2
63X=10
X=10/63
PROOF:
5*10/63-(4/5*10/63)=2/3
50/63-4*10/5*63)=2/3
50/63-40/315=2/3
50/63-8/63=2/3
42/63=2/3
2/3=2/3